OpenStudy (lena772):

Determine whether the points are co-linear. (Three points are co-linear if they lie on the same line.) (0, 4) (7, -6) (-5, 11)

OpenStudy (anonymous):

three pionts like (x1,y1),(x2,y2)& (x3,y3) are said to be co-linear iff: 1/2*{x1(y2-y3)+x2(y3-y1)+x3(y1-y2)}=0

OpenStudy (ddcamp):

Assume they are co-linear. Then they all lie on some line y = mx + b From point 1 (the y-intercept) we can see that b = 4 (4 = m(0) + b) Plugging point 2 into this, we have: -6 = 7m + 4 -10 = 7m m = -10/7 Giving us a final equation: y = (-10/7)x + 4 Does the third point fit this equation?

OpenStudy (lena772):

11= -5m +4 7=-5m m=-7/5

OpenStudy (lena772):

It doesn't fit right? @DDCamp

OpenStudy (ddcamp):

Right.

OpenStudy (lena772):

That means they're not co-linear?

OpenStudy (ddcamp):

Yes, they are not co-linear